# On structural contraction of biological interaction networks

### v3

@@ 2.4.1 Concentration dynamics as a quotient of extent-of-reaction dynamics
Fix a vector ${\bar{x}}\in\mathbb{R}_{+}^{n}$. Consider a system $\dot{\xi}=g(\xi)=R({\bar{x}}+\Gamma x)$, $\xi\in\mathbb{R}^{\nu}$, and let $\psi_{t}:\mathbb{R}^{\nu}\to\mathbb{R}^{\nu}$ be the associated flow. Similarly, consider the system $\dot{x}=f(x)=\Gamma R(x)$, $\xi\in\mathscr{C}_{\bar{x}}:={\bar{x}}+{\mathop{\rm Im\,}}\Gamma$, and let $\varphi_{t}:\mathscr{C}_{\bar{x}}\to\mathscr{C}_{\bar{x}}$ be the associated flow. Define the following mapping $\Phi:\mathbb{R}^{\nu}\to\mathscr{C}_{\bar{x}}$:
@@
Given a network ${\mathscr{N}}$ and a $\Gamma$-invariant GLF ${\widetilde{V}}$. Let $G:{\mathop{\rm Im\,}}\Gamma\to\mathbb{R}^{\nu}\!/\!{\mathop{\rm Ker\,}}\Gamma$ be as defined above. Then, ${\widehat{V}}:={\widetilde{V}}\circ G$ is called a dual GLF.
Using Theorem 2.7, we will shortly show in §2.4.3 that $({\widetilde{V}}\circ\Psi)(x)={\widehat{V}}(x-{\bar{x}})$ is also a Lyapunov function on $\mathscr{C}_{\bar{x}}$. However, we will first establish that ${\widetilde{V}}$ as a GLF for the extent-of-reaction dynamics, in a sense analogous to GLF’s defined earlier.
@@ 2.4.2 A GLF for the extent-of-reaction dynamics
Consider a network ${\mathscr{N}}=({\mathscr{S}},{\mathscr{R}})$, and let $R$ be an admissible rate. Consider the extent-of-reaction system (8) $\dot{\xi}=R({\bar{x}}+\Gamma\xi)$. Assume that ${\bar{x}}$ is a steady state for (4). If ${\widetilde{V}}$ is a $\Gamma$-translation invariant GLF for ${\mathscr{N}}$, then ${\widetilde{V}}$ a Lyapunov function for the extent-of-reaction system (8).
@@
We show the four statements given in Definition 2.9. The first and second statements follow immediately from the corresponding properties of ${\widetilde{V}}$ given in Defintion 2.1. The third statement is assumed in the statement of the proposition. We prove the fourth statement next.
Since ${\widetilde{V}}$ is a locally Lipschitz function, its gradient exists almost everywhere. At the points where it is defined, one write $\dot{\widetilde{V}}(\xi)$ as follows:
@@
[…]
where the last inequality follows since ${\widetilde{V}}$ is a common Lyapunov function for the systems $\dot{r}_{1}=Q_{1}r,...,\dot{r}_{s}=Q_{s}r$. Since ${\widetilde{V}}$ is locally Lipschitz, the inequality holds for almost all $\xi$ using Radmacher’s theorem. Next, we interpret the gradient above in the sense of the upper Dini’s derivative (see Remark 2.2). Then we may use Lemma 9 in [11], which says that if the inequality above holds almost everywhere, then it holds for all $\xi\geq 0$. This completes the proof.
@@ 2.4.3 The dual GLF for concentration dynamics
Consider a network ${\mathscr{N}}$ with a steady state ${\bar{x}}$. Let $V={\widetilde{V}}\circ R$ be defined as in Proposition 2.5. Then, we claim that $V^{*}(x)={\widehat{V}}(x-{\bar{x}})$ is a Lyapunov function for (4) whenever $x(0)\in\mathscr{C}_{\bar{x}}$. Positive definiteness is evident. We can verifying the non-increasing property of ${\widetilde{V}}$ by writing:
@@
    If ${\widehat{V}}$ is a dual GLF for ${\mathscr{N}}$, then ${\widetilde{V}}:={\widehat{V}}\circ\Gamma$ is a $\Gamma$-translation-invariant GLF for ${\mathscr{N}}$.
@@
Consider a network ${\mathscr{N}}=({\mathscr{S}},{\mathscr{R}})$, and consider (4) with a fixed admissible $R$ and a steady state ${\bar{x}}$. Assume that ${\widehat{V}}$ is a dual GLF for ${\mathscr{N}}$. Then, $V(x)={\widehat{V}}(x-{\bar{x}}),x\in\mathscr{C}_{\bar{x}}$ is Lyapunov function for (4) on $\mathscr{C}_{\bar{x}}$.
@@ 2.5 The class of $\ell_{\infty}$-norm GLFs
We focus now on GLFs of the form ${\widetilde{V}}(r)=\|Cr\|_{\infty}$ for some matrix $C=[c_{1},..,c_{m}]^{T}\in\mathbb{R}^{m\times\nu}$, $m>0$. We will assume, without loss of generality, that $C$ is a minimal representation of ${\widetilde{V}}$ which means that for all $k=1,..,m$, the set $\{r|{\widetilde{V}}(r)=c_{k}^{T}r\}$ has non-empty interior. We can make this assumption because if $C$ is not minimal, we can remove rows from $C$ to make it a minimal representation.
@@
    ${\mathop{\rm Ker\,}}C={\mathop{\rm Ker\,}}\Gamma$ and there exist Metzler matrices $\tilde{\Lambda}_{\ell}\in\mathbb{R}^{2m\times 2m},\ell=1,..,s$ such that
    (16) $\tilde{C}Q_{\ell}=\tilde{\Lambda}_{\ell}\tilde{C},~{}\mbox{and}~{}\tilde{\Lambda}_{\ell}\mathbf{1}=0,\ell=1,..,s,$
@@
  • There exists a matrix $B$ with $\mathop{\mbox{rank}}\Gamma=\mathop{\mbox{rank}}(B\Gamma)$ such that $C=B\Gamma$ and there exist matrices $\Lambda_{\ell}\in\mathbb{R}^{m\times m},\ell=1,..,s$ such that
    $CQ_{\ell}=\Lambda_{\ell}C~{}\mbox{and}~{}\mu_{\infty}(\Lambda_{\ell})\leq 0,\ell=1,..,s.$
@@
Using Corollary 2.12, and recalling that $C$ can be written as $C=B\Gamma$, we get the function ${\widehat{V}}$ as:
@@
Consider the BIN (9) and the introduced GLF (11). Corollary 2.15 asserts the existence of six matrix $\Lambda_{\ell},\ell=1,..,s$. Using the algorithms developed in [4, 5] we can find the matrices as follows:
@@
$\Lambda_{4}$ $=\left[\begin{array}[]{rrrrrrrr}0&0&0&0&0&0\\ 0&0&0&0&0&0\\ 0&-1&-1&0&0&0\\ 0&0&0&-1&0&0\\ 0&0&0&0&0&0\\ 0&0&0&0&1&-1\end{array}\right],\Lambda_{5}=\left[\begin{array}[]{rrrrrrrr}0&0&0&0&0&0\\ 0&0&0&0&0&0\\ -1&0&-1&0&0&0\\ 0&0&0&-1&1&0\\ 0&0&0&0&0&0\\ 0&0&0&0&0&-1\end{array}\right],\Lambda_{6}=\left[\begin{array}[]{rrrrrrrr}-1&0&-1&0&0&0\\ 0&0&0&0&0&0\\ 0&0&0&0&0&0\\ 0&0&0&0&0&0\\ 0&0&0&1&-1&0\\ 0&0&0&0&0&-1\end{array}\right]$
@@ 3.1 Preliminaries
The are several formulations of contraction theory. We are going to present the formulation that utilizes logarithmic norms (a.k.a, matrix measures).
@@
Let $(\mathbb{R}^{,}|.|)$ be a normed space, and let $\|.\|_{*}$ be the corresponding induced matrix norm on $\mathbb{R}^{n\times n}$. Then, the associated matrix measure (a.k.a, logarithmic norm) can be defined as follows for a matrix $A\in\mathbb{R}^{n\times n}$:
(18) $\mu_{*}(A):=\lim_{h\to 0^{+}}\frac{\|I+hA\|_{*}-1}{h},$
@@ 3.2 Nonexpansion in concentration coordinates
Our approach to prove the theorem is to show that the appropriate logarithmic norm is non-positive. However, since Theorem 3.4 is stated with respect to the functional |.$|.|_{B}$ on the vector space ${\mathop{\rm Im\,}}\Gamma$, we need to introduce the necessary notions.
@@ 3.2.1 Norms with non-square weightings
Let $(\mathbb{R}^{n},|.|)$ be a normed space. Let $B\in\mathbb{R}^{m\times n}$. We are interested in studying the pair $({\mathcal{V}},|.|_{B})$ where ${\mathcal{V}}$ is a vector subspace of $\mathbb{R}^{n}$ and |.$|.|_{B}$ is the weighted operator |.$|.|_{B}:z\mapsto|Bz|$. Such operators are usually studied using the notion of Minkowski functionals or gauge functions [80]. However, to keep the discussion self-contained, we develop the required notions here. For easier future reference, we state a simple equivalency first:
@@
    ${\mathop{\rm Im\,}}(B\Gamma)={\mathop{\rm Im\,}}(\Gamma)$.
@@
Let $(\mathbb{R}^{n},|.|)$ be a normed space. Let $B\in\mathbb{R}^{m\times n}$, and let ${\mathcal{V}}\subset\mathbb{R}^{n}$ be a vector subspace. If $B{\mathcal{V}}={\mathcal{V}}$, then the pair $({\mathcal{V}},|.|_{B})$ is a normed space.
@@
Since ${\mathcal{V}}$ is a vector subspace of $\mathbb{R}^{n}$, we immediately get that |.$|.|_{B}$ is semi-norm on ${\mathcal{V}}$. To show that it is a norm, we need to show that for all $x\in{\mathcal{V}}$, $|x|_{B}=0$ iff $x=0$. But this follows immediately from Lemma 3.5 since we assumed that $B{\mathcal{V}}={\mathcal{V}}$
@@ 3.2.2 Induced matrix norms and logarithmic norms restricted to a subspace
Similarly, we need to define the corresponding induced $(B,{\mathcal{V}})$-matrix norm of a matrix $J\in\mathbb{R}^{n\times n}$ as we did above. As in Lemma 3.6, we assume that $B{\mathcal{V}}={\mathcal{V}}$, i.e., the dimension of ${\mathcal{V}}$ does not reduce if we multiply each vector by $B$ from the left. Using Lemma 3.5, any vector $z\in{\mathcal{V}}$ can be also written as $LBz$ for some matrix $L\in\mathbb{R}^{m\times n}$. Therefore, we can write:
@@
Consider the normed space $(\mathbb{R}^{n},|.|)$. Let ${\mathcal{V}}\subset\mathbb{R}^{n}$ be a vector space, and let $B\in\mathbb{R}^{m\times n}$ be a matrix satisfying $B{\mathcal{V}}={\mathcal{V}}$. Then,
@@ 3.2.3 Computation of the logarithmic norm
Let a network ${\mathscr{N}}=({\mathscr{R}},{\mathscr{S}})$ be given. Fix an arbitrary ${\bar{x}}\in\mathbb{R}_{\geq 0}^{n}$. Assume that there exists an GLF for ${\mathscr{N}}$ of the form ${\widetilde{V}}(r)=\|B\Gamma r\|_{\infty}$. Then,
@@
where $\rho_{1},..,\rho_{s}$ are the positive elements in $K$ as in (1), and $\Lambda_{1},..,\Lambda_{s}$ are constructed by solving the convex program in Corollary 2.15.
@@
(21) $\Gamma K=\sum_{\ell=1}^{s}\bar{\rho}_{\ell}\Gamma E_{j_{\ell}i_{\ell}}=\sum_{\ell=1}^{s}\bar{\rho}_{\ell}\overbrace{(\Gamma_{j_{\ell}}e_{i_{\ell}}^{T})}^{J_{\ell}}=\sum_{\ell=1}^{s}\bar{\rho}_{\ell}J_{\ell},$
where $\Gamma_{1},..,\Gamma_{n}$ are the columns of $\Gamma$. In addition, we need the following basic lemma:
@@
Given a network ${\mathscr{N}}$. Let $J_{\ell},Q_{\ell},\ell_{1},..,s$ be defined as above. Then:
@@
    Let $C,B,\Lambda_{\ell},\ell=1,..,s$ be defined as in Corollary 2.15, and let $D$ be matrix whose rows form a basis for the left null space of $\Gamma$. then there exist matrices $Y_{1},..,Y_{s}$ such that $BJ_{\ell}=\Lambda_{\ell}B+Y_{\ell}D$, $\ell=1,..,s$.
@@
Therefore, the rows of $BJ_{\ell}-\Lambda_{\ell}B$ belong to the left null space of $\Gamma$. Hence, there exist matrices $Y_{1},..,Y_{s}$ such that $BJ_{\ell}-\Lambda_{\ell}B=Y_{\ell}D$, $\ell=1,..,s$. By rearranging we get the required expression.
@@
where the equality ()($\star$)( ⋆ ) follows from (21), equality ()($\clubsuit$)( ♣ ) follows from Lemma 3.10, and equality ()($\spadesuit$)( ♠ ) follows since $Dz=0$ for all $z\in{\mathop{\rm Im\,}}\Gamma$.

### v4

@@ 2.4.1 Concentration dynamics as a quotient of extent-of-reaction dynamics
Fix a vector ${\bar{x}}\in\mathbb{R}_{+}^{n}$. Consider a system $\dot{\xi}=g(\xi)=R({\bar{x}}+\Gamma{\xi})$, $\xi\in\mathbb{R}^{\nu}$, and let $\psi_{t}:\mathbb{R}^{\nu}\to\mathbb{R}^{\nu}$ be the associated flow. Similarly, consider the system $\dot{x}=f(x)=\Gamma R(x)$, $\xi\in\mathscr{C}_{\bar{x}}:={\bar{x}}+{\mathop{\rm Im\,}}\Gamma$, and let $\varphi_{t}:\mathscr{C}_{\bar{x}}\to\mathscr{C}_{\bar{x}}$ be the associated flow. Define the following mapping $\Phi:\mathbb{R}^{\nu}\to\mathscr{C}_{\bar{x}}$:
@@
Given a network $\mathscr{N}$} and a $\Gamma$-translation-invariant GLF $\widetilde{V}$}. Let $G:{\mathop{\rm Im\,}}\Gamma\to\mathbb{R}^{\nu}\!/\!{\mathop{\rm Ker\,}}\Gamma$ be as defined above. Then, ${\widehat{V}}:={\widetilde{V}}\circ G$ is called a dual GLF.
Using Theorem 2.7, we will shortly show in §2.4.3 that $({\widetilde{V}}\circ\Psi)(x)={\widehat{V}}(x-{\bar{x}})$ is also a Lyapunov function on $\mathscr{C}_{\bar{x}}$. However, we will first establish that $\widetilde{V}$} is a GLF for the extent-of-reaction dynamics, in a sense analogous to GLF’s defined earlier.
@@ 2.4.2 A GLF for the extent-of-reaction dynamics
Consider a network ${\mathscr{N}}=({\mathscr{S}},{\mathscr{R}})$, and let $R$ be an admissible rate. Consider the extent-of-reaction system (8) $\dot{\xi}=R({\bar{x}}+\Gamma\xi)$. Assume that $\bar{x}$} is a steady state for (4). If $\widetilde{V}$} is a $\Gamma$-translation-invariant GLF for $\mathscr{N}$}, then $\widetilde{V}$} is a Lyapunov function for the extent-of-reaction system (8).
@@
We show the four statements given in Definition 2.9. The first and second statements follow immediately from the corresponding properties of $\widetilde{V}$} given in Definition 2.1. The third statement is assumed in the statement of the proposition. We prove the fourth statement next.
Since $\widetilde{V}$} is a locally Lipschitz function, its gradient exists almost everywhere. At the points where it is defined, one writes $\dot{\widetilde{V}}(\xi)$ as follows:
@@
where $x_{\xi}^{*}:={\bar{x}}+\varepsilon\Gamma\xi$ for some $\varepsilon\in(0,1)$. The equality $(\star)$ above follows from applying the Mean-Value Theorem to the single-valued function $h(\varepsilon)=R({\bar{x}}+\varepsilon\Gamma\xi)$. The equality $(\star\star)$ can be shown as follows. Since $\widetilde{V}$} is $\Gamma$-translation-invariant, pick a $^{\widehat{V}}$ such that ${\widehat{V}}={\widetilde{V}}\circ G$ (see Theorem 2.7 and Definition 2.8). Hence, we can write ${\widetilde{V}}={\widehat{V}}\circ\Gamma$. Therefore, we can write $\partial{\widetilde{V}}/\partial\xi=(\partial{\widehat{V}}/\partial z)\Gamma$, but since $\bar{x}$} is a steady state, then $\Gamma R({\bar{x}})=0$, and hence the first term vanishes. Similar to (21), we can decompose the Jacobian as follows:
$\dot{\widetilde{V}}(\xi)\;=\;\sum_{\ell=1}^{s}\rho_{\ell}(x_{\xi}^{*})\frac{\partial{\widetilde{V}}}{\partial\xi}Q_{\ell}\xi\;\leq\;0\,$,
where the last inequality follows since $\widetilde{V}$} is a common Lyapunov function for the systems ${\dot{r}}=Q_{1}r,\dots,\dot{r}=Q_{s}r$. Since $\widetilde{V}$} is locally Lipschitz, the inequality holds for almost all $\xi$ using Radmacher’s theorem. Next, we interpret the gradient above in the sense of the upper Dini’s derivative (see Remark 2.2). Then we may use Lemma 9 in [11], which says that if the inequality above holds almost everywhere, then it holds for all $\xi\geq 0$. This completes the proof.
@@ 2.4.3 The dual GLF for concentration dynamics
As we will see in section 3, nonexpansivity of the trajectories in concentration coordinates can be understood with respect to the dual GLF defined on concentration coordinates, which we will explain next. Consider a network $\mathscr{N}$} with a steady state $\bar{x}$}. Let $V={\widetilde{V}}\circ R$ be defined as in Proposition 2.5. Then, we claim that $V^{*}(x)={\widehat{V}}(x-{\bar{x}})$ is a Lyapunov function for (4) whenever $x(0)\in\mathscr{C}_{\bar{x}}$. Positive definiteness is evident. We can verify the non-increasing property of $\widetilde{V}$} by writing:
@@
    If $^{\widehat{V}}$ is a dual GLF for $\mathscr{N}$}, then ${\widetilde{V}}:={\widehat{V}}\circ\Gamma$ is a $\Gamma$-translation-invariant GLF for $\mathscr{N}$}.
@@
Consider a network ${\mathscr{N}}=({\mathscr{S}},{\mathscr{R}})$, and consider (4) with a fixed admissible $R$ and a steady state $\bar{x}$}. Assume that $^{\widehat{V}}$ is a dual GLF for $\mathscr{N}$}. Then, $V(x)={\widehat{V}}(x-{\bar{x}}),x\in\mathscr{C}_{\bar{x}}$ is a Lyapunov function for (4) on $\mathscr{C}_{\bar{x}}$.
@@ 2.5 The class of $\ell_{\infty}$-norm GLFs
We focus now on GLFs of the form ${\widetilde{V}}(r)=\|Cr\|_{\infty}$ for some matrix $C=[c_{1},\dots,c_{m}]^{T}\in\mathbb{R}^{m\times\nu}$, $m>0$. We will assume, without loss of generality, that $C$ is a minimal representation of $\widetilde{V}$} which means that for all $k=1,\dots,m$, the set $\{r|{\widetilde{V}}(r)=c_{k}^{T}r\}$ has non-empty interior. We can make this assumption because if $C$ is not minimal, we can remove rows from $C$ to make it a minimal representation.
@@
    ${\mathop{\rm Ker\,}}C={\mathop{\rm Ker\,}}\Gamma$ and there exist Metzler matrices $\tilde{\Lambda}_{\ell}\in\mathbb{R}^{2m\times 2m},\ell=1,\dots,s$ such that
    (16) $\tilde{C}Q_{\ell}=\tilde{\Lambda}_{\ell}\tilde{C},~\mbox{and}~\tilde{\Lambda}_{\ell}\mathbf{1}=0,\ell=1,\dots,s,$
@@
  • There exists a matrix $B$ with $\mathop{\mbox{rank}}\Gamma=\mathop{\mbox{rank}}(B\Gamma)$ such that $C=B\Gamma$ and there exist matrices $\Lambda_{\ell}\in\mathbb{R}^{m\times m},\ell=1,\dots,s$ such that
    $CQ_{\ell}=\Lambda_{\ell}C~\mbox{and}~\mu_{\infty}(\Lambda_{\ell})\leq 0,\ell=1,\dots,s.$
@@
Using Corollary 2.12, and recalling that $C$ can be written as $C=B\Gamma$, we get the function $^{\widehat{V}}$ as:
@@
Consider the BIN (9) and the introduced GLF (11). Corollary 2.15 asserts the existence of six matrices $\Lambda_{\ell},\ell=1,\dots,s$. Using the algorithms developed in [4, 5] we can find the matrices as follows:
@@
$\Lambda_{4}$ $=\left[\begin{array}[]{rrrrrrrr}0&0&0&0&0&0\\ 0&0&0&0&0&0\\ 0&-1&-1&0&0&0\\ 0&0&0&-1&0&0\\ 0&0&0&0&0&0\\ 0&0&0&0&1&-1\end{array}\right],\Lambda_{5}=\left[\begin{array}[]{rrrrrrrr}0&0&0&0&0&0\\ 0&0&0&0&0&0\\ -1&0&-1&0&0&0\\ 0&0&0&-1&1&0\\ 0&0&0&0&0&0\\ 0&0&0&0&0&-1\end{array}\right],\Lambda_{6}=\left[\begin{array}[]{rrrrrrrr}-1&0&-1&0&0&0\\ 0&0&0&0&0&0\\ 0&0&0&0&0&0\\ 0&0&0&0&0&0\\ 0&0&0&1&-1&0\\ 0&0&0&0&0&-1\end{array}\right].$
@@ 3.1 Preliminaries
The are several formulations of contraction theory. We are going to present the formulation that utilizes logarithmic norms (a.k.a. matrix measures).
@@
Let $(\mathbb{R}^{n},|.|)$ be a normed space, and let $\|.\|_{*}$ be the corresponding induced matrix norm on $\mathbb{R}^{n\times n}$. Then, the associated matrix measure (a.k.a. logarithmic norm) can be defined as follows for a matrix $A\in\mathbb{R}^{n\times n}$:
(18) $\mu_{*}(A):=\lim_{h\to 0^{+}}\frac{\|I+hA\|_{*}-1}{h}{.}$
@@ 3.2 Nonexpansion in concentration coordinates
Although the theorem assumes the existence of a GLF defined on reaction coordinates, the trajectories are nonexpansive with respect to the norm $|.|_{B}$ which parallels the structure of the dual GLF defined earlier in §2. Our approach to prove the theorem is to show that the appropriate logarithmic norm is non-positive. However, since Theorem 3.4 is stated with respect to the functional $|.|_{B}$ on the vector space ${\mathop{\rm Im\,}}\Gamma$, we need to introduce the necessary notions.
@@ 3.2.1 Norms with non-square weightings
Consider normed spaces of the form $(\mathbb{R}^{n},|.|),n\geq 1$. Let $B\in\mathbb{R}^{m\times n}$. We are interested in studying the pair $({\mathcal{V}},|.|_{B})$ where $\mathcal{V}$} is a vector subspace of $\mathbb{R}^{n}$ and $|.|_{B}$ is the weighted operator $|.|_{B}:z\mapsto|Bz|$, whose domain is $\mathbb{{\mathcal{V}}}\subset\mathbb{R}^{n}$. Such operators are usually studied using the notion of Minkowski functionals or gauge functions [80]. However, to keep the discussion self-contained, we develop the required notions here.
For easier future reference, we state a simple equivalency first:
@@
    ${\dim}({\mathop{\rm Im\,}}(B\Gamma))={\dim}({\mathop{\rm Im\,}}(\Gamma))$.
@@
Let $(\mathbb{R}^{n},|.|)$ be a normed space. Let $B\in\mathbb{R}^{m\times n}$, and let ${\mathcal{V}}\subset\mathbb{R}^{n}$ be a vector subspace. If ${\dim}(B{\mathcal{V}})={\dim}({\mathcal{V}})$, then the pair $({\mathcal{V}},|.|_{B})$ is a normed space.
@@
Since $\mathcal{V}$} is a vector subspace of $\mathbb{R}^{n}$, we immediately get that $|.|_{B}$ is semi-norm on $\mathcal{V}$}. To show that it is a norm, we need to show that for all $x\in{\mathcal{V}}$, $|x|_{B}=0$ iff $x=0$. But this follows immediately from Lemma 3.5 since we assumed that ${\dim}(B{\mathcal{V}})={\dim}({\mathcal{V}})$.
@@ 3.2.2 Induced matrix norms and logarithmic norms restricted to a subspace
Similarly, we need to define the corresponding induced $(B,{\mathcal{V}})$-matrix norm of a matrix $J\in\mathbb{R}^{n\times n}$ as we did above. As in Lemma 3.6, we assume that ${\dim}(B{\mathcal{V}})={\dim}({\mathcal{V}})$, i.e., the dimension of $\mathcal{V}$} does not reduce if we multiply each vector by $B$ from the left. Using Lemma 3.5, any vector $z\in{\mathcal{V}}$ can be also written as $LBz$ for some matrix $L\in\mathbb{R}^{{n\times m}}$. Therefore, we can write:
@@
Consider the normed space $(\mathbb{R}^{n},|.|)$. Let ${\mathcal{V}}\subset\mathbb{R}^{n}$ be a vector space, and let $B\in\mathbb{R}^{m\times n}$ be a matrix satisfying ${\dim}(B{\mathcal{V}})={\dim}({\mathcal{V}})$. Then,
@@ 3.2.3 Computation of the logarithmic norm
Let a network ${\mathscr{N}}=({\mathscr{R}},{\mathscr{S}})$ be given. Fix an arbitrary ${\bar{x}}\in\mathbb{R}_{\geq 0}^{n}$. Assume that there exists a GLF for $\mathscr{N}$} of the form ${\widetilde{V}}(r)=\|B\Gamma r\|_{\infty}$. Then,
@@
where $\rho_{1},\dots,\rho_{s}$ are the positive elements in $K$ as in (1), and $\Lambda_{1},\dots,\Lambda_{s}$ are constructed by solving the convex program in Corollary 2.15.
@@
(21) $\Gamma K=\sum_{\ell=1}^{s}\bar{\rho}_{\ell}\Gamma E_{j_{\ell}i_{\ell}}={\sum_{\ell=1}^{s}\bar{\rho}_{\ell}\Gamma e_{j_{\ell}}e_{i_{\ell}}^{T}}=\sum_{\ell=1}^{s}\bar{\rho}_{\ell}\overbrace{(\Gamma_{j_{\ell}}e_{i_{\ell}}^{T})}^{J_{\ell}}=\sum_{\ell=1}^{s}\bar{\rho}_{\ell}J_{\ell},$
where $\Gamma_{1},\dots,\Gamma_{n}$ are the columns of $\Gamma$. In addition, we need the following basic lemma:
@@
Given a network $\mathscr{N}$}. Let $J_{\ell},Q_{\ell},{\ell=1,\ldots,s}$ be defined as above. Then:
@@
    Let $C,B,\Lambda_{\ell},\ell=1,\dots,s$ be defined as in Corollary 2.15, and let $D$ be a matrix whose rows form a basis for the left null space of $\Gamma$. Then there exist matrices $Y_{1},\dots,Y_{s}$ such that $BJ_{\ell}=\Lambda_{\ell}B+Y_{\ell}D$, $\ell=1,{\ldots},s$.
@@
Therefore, the rows of $BJ_{\ell}-\Lambda_{\ell}B$ belong to the left null space of $\Gamma$. Hence, there exist matrices $Y_{1},\dots,Y_{s}$ such that $BJ_{\ell}-\Lambda_{\ell}B=Y_{\ell}D$, $\ell=1,\dots,s$. By rearranging we get the required expression.
@@
where the equality $(\star)$ follows from (21), equality $(\clubsuit)$ follows from Lemma 3.10, and equality $(\spadesuit)$ follows since $Dz=0$ for all $z\in{\mathop{\rm Im\,}}\Gamma$.